I am using a circle graph it says 200 people are employed in a department store it's asking how many employees are in each section like women's department store they're are 25% employees how do I figure it out do I divide 200 by 25% can you help me
Answer by KMST(5347)

25% means 25 people out of a group of 100 people.
Put two such groups together and you have 100+100=200 people total,
with 25+25=50 in the women's section.
In general, you can view percentages as ratios or fractions.
25% = 25/100 =1/4 = 0.25
Whatever the total is, to find how much of that total is a certain percentage
you just multiply the total times the fraction or decimal form of the percentage.
For example,
37% of 146 people is 146 × 0.37 , and
125% of $38,000 is $38,000 × 1.25 .